'n' no of batteries of emf 'e' and internal resistance 'r' are available. How do you arrange the batteries to obtain the maximum current in external resistance 'R'. Find the maximum current in external resistance.

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Greatest current in outside resistance can be accomplished by organizing every one of the batteries in arrangement. 
In arrangement condition, 
net emf = ne 
net resistance=r/n 
current through R=ne/(R+r/n) 
since r is little can be disregarded. 
current = ne/R 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)